# One-Shot Imitation under Mismatched Execution

### v1

@@ Abstract
Human demonstrations as prompts are a powerful way to program robots to do long-horizon manipulation tasks. However, directly translating such demonstrations into robot-executable actions poses significant challenges due to execution mismatches, such as different movement styles and physical capabilities. Existing methods either rely on robot-demonstrator paired data, which is infeasible to scale, or overly rely on frame-level visual similarities, which fail to hold. To address these challenges, we propose RHyME, a novel framework that automatically establishes task execution correspondences between the robot and the demonstrator by using optimal transport costs. Given long-horizon robot demonstrations, RHyME synthesizes semantically equivalent human demonstrations by retrieving and composing similar short-horizon human clips, facilitating effective policy training without the need for paired data. We show that RHyME outperforms a range of baselines across various cross-embodiment datasets on all degrees of mismatches. Through detailed analysis, we uncover insights for learning and leveraging cross-embodiment visual representations.
@@ I Introduction
Human demonstrations as prompts are a powerful way to program robots to do long-horizon manipulation tasks. Compared to language instructions, demonstrations are grounded in the task environment and are rich with cues that detail what steps to follow, which objects to interact with, and how to interact with them.
@@
We can view this as a translation problem where a human video must be translated into a series of robot actions [1, 2]. However, in the most naive setting, training such policies requires paired data of human and robot demonstrations, which are infeasible to collect at a large scale for long-horizon tasks. While there exists internet-scale human demonstration videos (e.g. YouTube) and increasingly large robot datasets [3, 4], these datasets are unpaired and cannot directly be used to learn the translation.
@@
Prior works attempt to leverage unpaired human and robot demonstrations to learn an aligned visual representation that map both human and robot images to a common embedding [5, 6]. A policy is then trained to generate actions conditioned on robot video embeddings, and directly transferred at test time to work with embeddings from a human prompt video. However, a key assumption these works rely on is that the demonstrator and robot perform tasks with matched execution, i.e. the human executes tasks in a visually similar way to that of the robot (e.g. slowly moving one arm with a simple grasp). Oftentimes, this is not the case: it is natural for humans to act more swiftly, use both hands to manipulate objects, or even execute two separate tasks simultaneously. This results in a misalignment between demonstrator and robot embeddings, preventing a direct transfer of policies.
@@
We tackle this problem of imitation under mismatched execution. Our key insight is that while a robot and demonstrator may perform the same task in visually and physically different ways, we can establish a high-level equivalence by reasoning over the entire sequence of image embeddings they generate. We show that while individual image embeddings may appear different between robot and demonstrator, we can construct sequence-level similarity functions where the two are closer. Notably, we can do this without fine-tuning representations on explicitly paired data.
@@
We propose RHyME (Retrieval for Hybrid Imitation under Mismatched Execution), a framework which trains a robot policy to follow a long-horizon demonstration from a mismatched expert without access to paired robot-demonstration videos (Fig 1). First, RHyME defines a sequence-level similarity metric on robot and demonstrator embeddings as the optimal transport cost. Given a robot trajectory and a database of demonstrator data, RHyME imagines a long-horizon video by retrieving and composing short-horizon snippets of demonstrations similar to the robot video. To do so, RHyME uses an optimal transport cost as a similarity metric between robot and demonstrator sequences. RHyME then trains a policy in a hybrid fashion on both robot demonstrations and imagined demonstrations. Our contributions:
Figure 1: Overview of RHyME. We introduce RHyME, a hierarchical framework that trains a robot policy to mimic a long-horizon video from a demonstrator that exhibits mismatched task execution. We first learn unsupervised visual embeddings shared by a robot and demonstrator, then employ sequence-level similarity metrics to ”imagine” long-horizon demonstrations that are used at train time.
@@
    We propose RHyMe, a retrieval-based algorithm using optimal transport to construct synthetic videos of a mismatched demonstrator from a robot video to be used for policy training.
@@
    We conduct analysis of different retrieval functions on novel cross-embodiment datasets on the Franka Kitchen simulator that exhibit varying levels of visual and execution mismatches between the robot and demonstrator.
@@
    We show that RHyMe outperforms a range of baselines on all degrees of mismatches and uncover insights for learning cross-embodiment visual representations.
@@ II Related Work
Tracking reference motion. The imitation challenge is reduced to motion tracking when the robot receives the demonstrator’s motion as input. Robots with human-like joint configurations can directly mimic human trajectories, a method applied to humanoid robots with similar morphologies [7, 8, 9, 10, 11, 12] and robotic hands [13, 14, 15, 16, 17, 18, 19, 20]. When the execution capabilities are mismatched, either human models are simplified [7, 9, 11] or robot trajectories [8, 10] are optimized to approximately match the reference trajectory. [12] shows a method to extract human poses from videos and use that as a reference for training reinforcement learning agents. Similarly, human demonstrations have been used to guide reinforcement learning for robotic hands [13, 14, 15, 16]. More recently, large-scale video datasets of humans on the internet have been used to extract hand positions and adapted for robot manipulation [19, 18, 20, 17]. Distinct from these works, we focus on learning robot manipulation tasks directly from demonstrators’ RGB videos without explicit motion input.
Learning reward functions from observations. These methods tackle the problem of matching the demonstrator when their motion cannot be simply mimicked. Still, their videos contain useful task information and can be used to learn reward functions for reinforcement learning on the robot. A general method in this line of research [21, 22, 23, 24] is to extract reward functions that encourage the robot to manipulate objects in the same way as the video. For example, [21] self-supervises the robot by making it learn to match the demonstrations of a demonstrator perturbing a rope. GraphIRL [24] extracts sequences of object pose movement from the demonstrator and enforces temporal cyclic consistency with the robot’s execution. Other approaches frame the problem as task matching [25, 26], where the robot is rewarded when it is deemed to perform the same skill as the demonstrator. While these papers tackle the problem of mismatched execution directly, they require reinforcement learning to train robot policies, which can be challenging for complex tasks.
@@
One-shot visual imitation from demonstration videos. We tackle this problem setting in our work where the robot imitates actions from demonstrator videos in a one-shot setting, i.e., the robot uses a prompt video as a guide, aiming to replicate the demonstrated actions after viewing them once [1, 2, 5, 6, 33, 34, 35]. If a paired dataset of robot and demonstrator videos executing the same task exists, the robot can learn to translate a prompt video into actions directly [1, 2]. The closest to our work is the setting without paired data of demonstrator and robot skills. Prior works [6, 5] train policies conditioning on robot videos and zero-shot transfer to a prompt demonstration at test time using aligned visual embeddings. For example, XSkill [5] uses a self-supervised clustering algorithm based on visual similarity to align representations of demonstrations and robot videos. However, such an approach can falter when there are significant mismatches in execution. We address this issue by posing the visual imitation problem as one of train-time retrieval. During training, we match robot videos to the closest demo videos from an unpaired play dataset to imagine synthetic demo videos. Training robot policies conditioned on these synthetic videos enable the robot to translate demonstration videos into robot actions at test time.
@@ III Problem Formulation
Test Time: Translate Demonstration Video to Robot Actions. The robot’s goal is to replicate a series of tasks as shown by the demonstrator using a policy $\pi(a_{R}|s_{R},\mathbf{v_{H}})$ that translates a demonstration video to predict the robot’s action $a_{R}$ at a robot state $s_{R}$. The demonstration video is a sequence of images $\mathbf{v_{H}}=\{v_{H}^{0},v_{H}^{1},\dots,v_{H}^{T}\}$, where $T$ is the length of the video.
@@
Train Time: Learning from Unpaired Demonstrator and Robot Data. Training the policy is straightforward given paired robot and demonstrator data. But, this is infeasible to collect at scale. Instead, we setup the problem as learning aligned embeddings from unpaired data that enable transferring policies trained on robot data embeddings to demonstration embeddings.
@@
Datasets. We assume access to two distinct datasets: a robot dataset of long-horizon manipulation tasks, denoted as $D_{robot}$, and a play dataset, denoted as $D_{play}$, consisting of short-horizon demonstrator (e.g., a human) video clips that show how to interact with different objects and the environment. The robot dataset, $D_{robot}=\{(\mathbf{\xi_{R}},\mathbf{v_{R}})\}$, comprises pairs of state-action trajectories and robot videos. Each robot trajectory, $\mathbf{\xi_{R}}=\{(s_{R}^{0},a_{R}^{0}),(s_{R}^{1},a_{R}^{1}),\dots,(s_{R}^{T},a_{R}^{T})\}$, represents the sequence of robot states and actions throughout an episode. Correspondingly, the video $\mathbf{v_{R}}=\{v_{R}^{0},v_{R}^{1},\dots,v_{R}^{T}\}$ is a sequence of images of the robot executing the task. The play dataset, $D_{play}=\{\mathbf{v_{H}}\}$, consists of demonstrator videos without any direct correspondence to the robot’s dataset. At test-time, the demonstrator’s video consists of a set of multiple tasks whose composition are never seen by the robot during training. However, we make an assumption that the constituent tasks are individually covered both in $D_{play}$ and $D_{robot}$. This assumption is consistent with prior work [5].
@@
Aligning Embeddings. To enable translation, we embed robot and demonstrator videos into a shared embedding space. We employ a video encoder $\phi(\mathbf{v})$ to extract a sequence of embeddings $\mathbf{z}=\{z_{0},z_{1},....,z_{T}\}$ corresponding to each frame from all videos [^1] We encode a 1-timestep sliding window of 8 neighboring images to generate each image embedding.. Then, following prior work [5], we train this encoder using an unsupervised loss function that clusters visually similar skills (see Appendix).
Transferring Policies. Using robot datapoints $(\mathbf{\xi_{R}},\mathbf{z_{R}})$, we first encode the robot video $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$, and then train a policy conditioned on these embeddings $\pi(a_{R}|s_{R},\mathbf{z_{R}})$. At test time, given a video from the demonstrator $\mathbf{v_{H}}$, we generate a latent embedding $\mathbf{z_{H}}$. Then, these embeddings are fed directly into the policy to predict actions $\pi(a_{R}|s_{R},\mathbf{z_{H}})$ that follow the demonstrator.
Problem of Mismatched Execution: The success of the approach hinges on the encoder’s ability to completely align individual robot and demonstrator embeddings. This may be infeasible due to significant discrepancies in execution.
@@ IV Approach
We present RHyME, a one-shot imitation learning algorithm that translates demonstrations to robot actions, without paired data. Before policy training, we first train a video encoder on using a dataset of unpaired robot and demonstration videos (Section  IV-A). Then, this trained video-encoder is frozen and utilized for retrieval during policy training (Section IV-B). At train time, given just a robot trajectory, RHyME imagines a corresponding demonstration by retrieving and composing corresponding demonstrator snippets. It then trains a policy to condition on this imagined demonstration to predict robot actions. We discuss details of the retrieval, training process, and video embeddings below.
IV-A Representation Learning of Visual Embeddings
We align embeddings of the robot and the demonstrator videos without trajectory-level correspondences in three different ways: visually, temporally and at the task-level. We employ unsupervised losses for visual and temporal alignment following prior works [36, 37] (see Appendix for details), and introduce an optional task alignment loss.
Figure 2: Performance on Mismatched Execution Datasets. We present results on three datasets (left). As the demonstrator’s actions visually and physically deviate further from those of the robot, policies trained with our framework RHyME consistently outperforms XSkill measured by task recall and imprecision rates.
Task Alignment. The task-level alignment is applied when there exists a small dataset of short-horizon snippets of the robot and demonstrator executing the same task. In contrast to the other frame-level representation learning methods used by prior methods, this method focuses on aligning video embeddings of the robot $\mathbf{z_{R}}$ and the demonstrator $\mathbf{z_{H}}$. To do so, we use the optimal transport distance function defined in Section IV-B, $d(\mathbf{z_{R}},\mathbf{z_{H}})$ to measures the matching between two sequence of video embeddings that may differ in length. Our alignment attempts to pull embedding sequences of paired human-demonstrator tasks closer to each other while pushing those of different tasks away from each other. We achieve this result by applying the INFO-NCE [38] contrastive learning objective similarly as before. In a batch $B$, consisting of paired demonstrator and robot video snippets $\{\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}\}_{i=1}^{B}$, we generate negative pairs by matching video embeddings from distinct pairs within the batch. Then, we construct the following loss function:
$L_{task}(\phi)=-\sum\limits_{i}\frac{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))}{\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{i}))+\sum_{j\neq i}\text{exp}(-d(\mathbf{z_{R}}^{i},\mathbf{z_{H}}^{j}))}$.
Input: Robot Embeddings $\mathbf{z_{R}}$, Demonstrator Play Dataset $D_{play}$, Video Encoder $\phi(\mathbf{z}|\mathbf{v})$, Segment Length $K$, Distance Function $d$
@@ IV-A Representation Learning of Visual Embeddings
// Divide long-horizon robot sequence into short-horizon clips
@@
// Find closest short-horizon clip embedding in play dataset
$\mathbf{\mathbf{\hat{z}_{play}}}\leftarrow\operatorname*{arg\,min}{d(\mathbf{z_{play}},{z_{R}^{i:i+K}}})$
// Extend imagined embedding sequence with retrieved demo
2      $\mathbf{\hat{z}_{H}}\text{.extend}(\mathbf{\hat{z}_{play}})$
Return Imagined Demo $\mathbf{\hat{z}_{H}}$
Algorithm 1 Imagine-Demo: Retrieving Matched Demonstrator Embeddings
Input: Robot Dataset ${D}_{robot}$, Demonstrator Play Dataset $D_{play}$, Video Encoder $\phi(\mathbf{z}|\mathbf{v})$
@@
Initialize Robot Policy $\pi_{\theta}$
1while not converged do
Get robot video and actions $\mathbf{\xi_{R}},\mathbf{v_{R}}\sim D_{robot}$
Generate robot embeddings $\mathbf{z_{R}}=\phi(v_{R})$
// Retrieve demonstrator embeddings from play data
$\mathbf{\hat{z}_{H}}\leftarrow\texttt{Imagine-Demo}(\mathbf{z_{R}},\mathbb{D}_{play})$

### v2

@@ Abstract
Human demonstrations as prompts are a powerful way to program robots to do long-horizon manipulation tasks. However, translating these demonstrations into robot-executable actions presents significant challenges due to execution mismatches in movement styles and physical capabilities. Existing methods either depend on robot-demonstrator paired data, which is infeasible to scale, or rely too heavily on frame-level visual similarities that often break down in practice. To address these challenges, we propose RHyME, a novel framework that automatically aligns robot and demonstrator task executions using optimal transport costs. Given long-horizon robot demonstrations, RHyME synthesizes semantically equivalent demonstrator videos by retrieving and composing short-horizon demonstrator clips. This approach facilitates effective policy training without the need for paired data. We demonstrate that RHyME outperforms a range of baselines across cross-embodiment datasets, showing a 52% increase in task recall over prior cross-embodiment learning methods. We release our code and datasets at website.
Figure 1: Overview of RHyME. We introduce RHyME, a hierarchical framework that trains a robot policy to mimic a long-horizon video from a demonstrator that exhibits mismatched task execution. Inference Time (Left): Our robot policy translates a demonstrator video into actions to complete the same long-horizon tasks specified by the input. Train Time (Right): Given unpaired robot-demonstrator datasets, RHyME ”imagines” a paired dataset by employing sequence-level similarity metrics which can be used for training the policy.
@@ I Introduction
Human demonstrations offer an effective approach for programming robots to execute long-horizon manipulation tasks [1, 2, 3, 4]. Unlike language instructions, demonstrations are grounded in the task environment, providing rich with cues for what steps to follow, which objects to interact with, and how to interact with them [5, 6].
@@
We view this as a translation problem where a human video must be translated into a series of robot actions [7, 8, 9, 10]. However, training such policies typically requires paired human-robot demonstrations, which is impractical to collect at scale for long-horizon tasks. Although large-scale human demonstration videos (e.g., YouTube) and robot datasets exist [11, 12], they are unpaired, making them unsuitable for directly learning this translation.
@@
Prior works leverage unpaired human and robot demonstrations to learn visual representations that map both human and robot images into a shared embedding space [4, 6, 13, 14, 15]. A policy is then trained to generate actions conditioned on robot video embeddings, and directly transferred at test time to work with embeddings from a human prompt video. However, a key assumption these works rely on is that the demonstrator and robot perform tasks with matched execution, i.e., the human executes tasks in a visually similar way to that of the robot (e.g. slowly moving one arm with a simple grasp). In reality, humans often act more swiftly, use both hands for manipulation, or even execute multiple tasks simultaneously, creating a mismatch in execution styles. This mismatch leads to misalignment between the demonstrator and robot embeddings, hindering direct policy transfer.
@@
We tackle this problem of imitation under mismatched execution. Our key insight is that while a robot and demonstrator may perform the same task in visually and physically different ways, we can establish a high-level equivalence by reasoning over the entire sequence of image embeddings they generate. We show that while individual image embeddings may appear different between robot and demonstrator, we can construct sequence-level similarity functions where the two are closer. Notably, we can do this without fine-tuning representations on paired data.
@@
We propose RHyME (Retrieval for Hybrid Imitation under Mismatched Execution), a framework which trains a robot policy to follow a long-horizon demonstration from a mismatched expert without access to paired robot-demonstration videos (Fig. 1). First, RHyME defines a sequence-level similarity metric between robot and demonstrator embeddings, using optimal transport to measure alignment. Given a robot trajectory and a database of demonstrator data, RHyME imagines a long-horizon video by retrieving and composing short-horizon snippets of demonstrations similar to the robot video. This retrieval process is guided by an optimal transport similarity metric between robot and demonstrator sequences. The framework trains a policy using a hybrid approach, incorporating both real robot demonstrations and imagined demonstrator sequences. Our contributions are:
@@
    We propose RHyME, a novel retrieval-based algorithm for one-shot initiation from videos of a demonstrator with mismatched execution. Without access to paired datasets, RHyME utilizes high-level equivalence between robot and demonstrator videos using optimal transport costs.
@@
    We systematically study the performance of different video retrieval by proposing three novel cross-embodiment datasets in the Franka Kitchen simulator. The demonstration datasets, totaling 10 hours of videos, are designed to vary in visual and execution mismatches relative to the robot.
@@
    We show that RHyMe outperforms a range of baselines on all degrees of mismatches, yielding over 50% increase in success rates in the most challenging scenarios.
@@ II Related Work
Tracking Reference Motion. The imitation challenge is reduced to motion tracking when the robot receives the demonstrator’s motion as input. Robots with human-like joint configurations can directly mimic human trajectories, a method applied to humanoid robots with similar morphologies [16, 17, 18, 19, 20, 21] and robotic hands [22, 23, 24, 25, 26, 27, 28, 29]. When the execution capabilities are mismatched, either human models are simplified [16, 18, 20] or robot trajectories [17, 19, 30, 31] are optimized to approximately match the reference trajectory. [21] shows a method to extract human poses from videos and use that as a reference for training reinforcement learning agents. Similarly, human demonstrations have been used to guide reinforcement learning for robotic hands [22, 23, 24, 25]. More recently, large-scale video datasets of humans on the internet have been used to extract hand positions and adapted for robot manipulation [28, 27, 29, 26, 32], or rely on other common abstractions such as optical flow as a common trajectory representation across embodiments [33], still requiring execution to be matched. Distinct from these works, we focus on learning robot manipulation tasks directly from demonstrators’ RGB videos without explicit motion input.
Learning Reward Functions from Demonstrator Videos. These methods tackle the problem of matching the demonstrator when their motion cannot be simply mimicked. Still, their videos contain useful task information and can be used to learn reward functions for reinforcement learning on the robot. A general method in this line of research [34, 35, 36, 2] is to extract reward functions that encourage the robot to manipulate objects in the same way as the video. For example, [34] self-supervises the robot by making it learn to match the demonstrations of a demonstrator perturbing a rope. GraphIRL [2] extracts sequences of object pose movement from the demonstrator and enforces temporal cyclic consistency with the robot’s execution. Other approaches frame the problem as task matching [37, 38], where the robot is rewarded when it is deemed to perform the same skill as the demonstrator. While these papers tackle the problem of mismatched execution directly, they require reinforcement learning to train robot policies, which can be challenging for complex tasks.
@@
One-shot Visual Imitation from Demonstration Videos. We tackle this problem setting in our work where the robot imitates actions from demonstrator videos in a one-shot setting, i.e., the robot uses a prompt video as a guide, aiming to replicate the demonstrated actions after viewing them once [7, 8, 4, 6, 41, 42, 43, 32]. If a paired dataset of robot and demonstrator videos executing the same task exists, the robot can learn to translate a prompt video into actions directly [7, 8]. The closest to our work is the setting without paired data of demonstrator and robot skills. Prior works [6, 4] train policies conditioning on robot videos and zero-shot transfer to a prompt demonstration at test time using aligned visual embeddings. For example, XSkill [4] uses a self-supervised clustering algorithm based on visual similarity to align representations of demonstrations and robot videos. However, such an approach can falter when there are significant mismatches in execution. We address this issue by posing the visual imitation problem as one of train-time retrieval. During training, we match robot videos to the closest demo videos from an unpaired play dataset to imagine synthetic demo videos. Training robot policies conditioned on these synthetic videos enable the robot to translate demonstration videos into robot actions at test time.
@@ III Problem Formulation
Inference Time: Translate Demonstration Video to Robot Actions. The robot’s goal is to replicate a series of tasks demonstrated in a video using a policy $\pi(a_{R}|s_{R},\mathbf{v_{H}})$ that translates the video into robot actions $a_{R}$ at state $s_{R}$. The demonstration video is a sequence of images $\mathbf{v_{H}}=\{v_{H}^{0},v_{H}^{1},\dots,v_{H}^{T}\}$, where $T$ is the length of the video.
@@
Train Time: Learning from Unpaired Demonstrator and Robot Data. While training a policy with paired robot and demonstrator data is feasible, collecting such data at scale is impractical. Instead, we frame the problem as learning and leveraging aligned embeddings from unpaired data, enabling the transfer of policies trained on robot embeddings to demonstration embeddings.
@@
We assume access to two datasets —a robot dataset ($D_{\rm robot}$) of long-horizon manipulation tasks and a play dataset ($D_{\rm play}$) of short-horizon demonstrator video clips showing interactions with objects and the environment. The robot dataset, $D_{\rm robot}=\{(\mathbf{\xi_{R}},\mathbf{v_{R}})\}$, comprises pairs of state-action trajectories and robot videos. Each robot trajectory, $\mathbf{\xi_{R}}=\{(s_{R}^{0},a_{R}^{0}),(s_{R}^{1},a_{R}^{1}),\dots,(s_{R}^{T},a_{R}^{T})\}$, represents the sequence of robot states and actions throughout an episode. Correspondingly, the video $\mathbf{v_{R}}=\{v_{R}^{0},v_{R}^{1},\dots,v_{R}^{T}\}$ is a sequence of images of the robot executing the task. The play dataset, $D_{\rm play}=\{\mathbf{v_{H}}\}$, consists of demonstrator videos that do not have direct correspondence with the robot dataset. At test time, the demonstrator’s video contains a set of tasks whose composition is unseen by the robot during training. However, we make an assumption that the constituent tasks are individually covered both in $D_{\rm play}$ and $D_{\rm robot}$. This assumption is consistent with prior work [4].
@@
Our goal is to train two modules: a vision encoder and a robot policy. The vision encoder maps both robot and demonstrator videos into a shared embedding space to enable translation. We employ a video encoder $\phi(\mathbf{v})$ to extract a sequence of embeddings $\mathbf{z}=\{z_{0},z_{1},....,z_{T}\}$ for each frame from all videos [^1] We encode a 1-timestep sliding window of 8 neighboring images to generate each image embedding.. Then, given a demonstrator video $\mathbf{v_{H}}$, we generate a sequence of latent embeddings $\mathbf{z_{H}}$. We aim to train a policy that conditions on the sequence of embeddings to predict robot actions $\pi(a_{R}|s_{R},\mathbf{z_{H}})$ without access to paired robot and demonstrator data. We discuss how to train both the encoder and the policy in Section IV.
@@ IV Approach
We present RHyME, a one-shot imitation learning algorithm that translates demonstrations to robot actions, without paired data. Before policy training, we first train a video encoder using a dataset of unpaired robot and demonstration videos (Section  IV-A). Then, this trained video-encoder is frozen and utilized for retrieval during policy training (Section IV-B). At train time, given just a robot trajectory, RHyME imagines a corresponding demonstration by retrieving and composing corresponding demonstrator snippets. It then trains a policy to predict robot actions, conditioned on the imagined demonstration. We discuss details of the retrieval, training process, and video embeddings below.
Input: Robot embeddings $\mathbf{z_{R}}$, demonstrator play dataset $D_{\rm play}$, video encoder $\phi(\mathbf{z}|\mathbf{v})$, segment length $K$, sequence-level video distance function $d$
@@
// Divide sequence into short snippets
@@
// Find closest snippet from play dataset
${{\hat{z}_{H}}}\leftarrow\operatorname*{arg\,min}\limits_{{\hat{z}_{H}}=\phi(v_{H}),v_{H}\in D_{\rm play}}{d({\hat{z}_{H}},{z_{R}^{i:i+K}}})$
// Append snippet to imagined demo
2      $\mathbf{\hat{z}_{H}}\text{.extend}({\hat{z}_{H}})$
Return Imagined demo embeddings $\mathbf{\hat{z}_{H}}$
Algorithm 1 Imagine-Demo: Retrieve and compose demonstration snippets given a robot trajectory
Input: Robot dataset ${D}_{\rm robot}$, demonstrator play dataset $D_{\rm play}$, video encoder $\phi(\mathbf{z}|\mathbf{v})$
@@
1Initialize robot policy $\pi_{\theta}$ and paired dataset $D_{\rm paired}\leftarrow\{\}$ for $(\mathbf{v_{R}},\mathbf{\xi_{R}})\text{ in }D_{\rm robot}$ do
Generate robot embeddings $\mathbf{z_{R}}=\phi(\mathbf{v_{R}})$
// Retrieve and pair demonstration
$\mathbf{\hat{z}_{H}}\leftarrow\texttt{Imagine-Demo}(\mathbf{z_{R}},{D}_{\rm play})$
// Add paired data into dataset
2      $D_{\rm paired}\leftarrow D_{\rm paired}\cup(\mathbf{\hat{z}_{H}},\mathbf{{z}_{R}},\mathbf{\xi_{R}})$
3while not converged do
// Get demo, robot embeddings with actions
$(\;\mathbf{\hat{z}_{H}},\mathbf{z_{R}},\mathbf{\xi_{R}}\;)\sim D_{\rm paired}$